What mass of carbon dioxide is present in 1.00 m3 of dry air at a temperature of 15  ∘C and a pressure of 689 torr?


1
Expert's answer
2021-02-24T03:42:08-0500

Pressure, "P" = 689 torr = 91859.1 Pa

Volume, "V" = 1 m3 = 1000 L

Temperature, "T" = "15^{\omicron}" C = 288 K

Molar mass of CO2, "M" = 44 g/mol

From Ideal gas equation,

"PV=nRT"

"\Rightarrow n=\dfrac{PV}{RT}"


"\Rightarrow\dfrac{m}{M}=\dfrac{PV}{RT}"


"\Rightarrow m=\dfrac{MPV}{RT}"


"\Rightarrow m=\dfrac{44\times91859.1\times1000}{8.314\times10^3\times288}"

"\Rightarrow m=1687.99\approx1688g"